# Trajectory Prediction for Autonomous Driving: Progress, Limitations, and Future Directions

### v1

@@ Abstract
As the potential for autonomous vehicles to be integrated on a large scale into modern traffic systems continues to grow, ensuring safe navigation in dynamic environments is crucial for smooth integration. To guarantee safety and prevent collisions, autonomous vehicles must be capable of accurately predicting the trajectories of surrounding traffic agents. Over the past decade, significant efforts from both academia and industry have been dedicated to designing solutions for precise trajectory forecasting. These efforts have produced a diverse range of approaches, raising questions about the differences between these methods and whether trajectory prediction challenges have been fully addressed. This paper reviews a substantial portion of recent trajectory prediction methods and devises a taxonomy to classify existing solutions. A general overview of the prediction pipeline is also provided, covering input and output modalities, modeling features, and prediction paradigms discussed in the literature. In addition, the paper discusses active research areas within trajectory prediction, addresses the posed research questions, and highlights the remaining research gaps and challenges.
@@ 1 Introduction
To accurately predict trajectories, the prediction model must account for the motion of vehicles, cyclists, pedestrians — all traffic participants. For vehicle motion, autonomous vehicles (AV) should be able to track surrounding vehicles and their potential maneuvers. As human drivers have an intuition of inter-dependency of surrounding vehicles, e.g., it can be anticipated if a neighboring car is cut off, there is a high chance that this vehicle will be compelled to perform a risky maneuver and potentially interfere with the ego-vehicle’s lane to avoid collision or suddenly hit the breaks. To equip autonomous vehicles with interaction-aware intuition, the prediction module needs to build a representation of each target agent’s surroundings and accounting for inter-dependency of traffic participants’ decisions. Regarding the human motion, beyond the complexity of interaction and inter-dependency, predicting human trajectories involves accounting for complex human behavior. Pedestrians’ motion is influenced not only by road topology and actions of surrounding agents but also by social rules, norms, and relations between agents. To accurately predict pedestrian trajectories, many factors need to be inferred from contextual and general knowledge information, e.g., a child and an adult exhibit different motion patterns. Additionally, prediction models that rely on traffic rules benefit from a reduced set of possible future trajectories. However, human drivers often violate traffic rules by over-speeding, changing lanes through a solid line, turning from a non-designated lane, ignoring stop signs, or even passing through a red light. Prediction models need to account for the risk of such violations to ensure robust and reliable predictions in real-world scenarios. Lastly, the trajectory prediction task has a multimodal nature, i.e., the same past trajectory can unfold into several possible future trajectories, e.g., a vehicle may slow down for a traffic signal or speed through if it turns green. To account for multimodality, the output of the prediction model must contain the distribution of possible trajectories with probabilities.
@@
To address these challenges, the research community has focused on designing models capable of capturing complex patterns essential for accurate prediction. Among the top-performing solutions is Trajectron++ [1], which integrates environmental context through a graph-structured recurrent architecture to model interactions between agents. The Spatio-Temporal Transformer Network (S2TNet) [2] leverages transformer layers to capture both spatial and temporal dependencies. Furthermore, from the intersection between academia and industry, the MultiPath++ model [3], released by the Waymo research group, employs multi-context gating fusion and dynamically learns reference points (anchors) to support goal-oriented predictions. Collectively, these models combine diverse techniques, such as graph-based interaction modeling, transformer architectures, and dynamic goal sampling, leading to more accurate and robust trajectory predictions. On the other side, the commercial research prioritizes real-time execution, cost-effectiveness, and handling uncertain situations. This leads each company to use its own combination of sensors, compatible with the set of algorithms it employs. Tesla’s prediction algorithm is an end-to-end vision-based deep learning model [^1] https://www.tesla.com/support/autopilot, forecasting trajectories based on the visual sequences and aggregated contextual information over time. The model continuously learns from the vast amount of data collected by Tesla’s global fleet of vehicles. In contrast, Waymo utilizes a combination of high-fidelity sensors (LiDAR, cameras, and radar). The sensor data is processed by the perception module and then passed to motion prediction and planning components. Waymo adopts a modular structure, ensuring that each component can be thoroughly tested independently and then integrated into a single pipeline [^2] https://waymo.com/research/. It is challenging to determine the specific algorithms employed in real operations, but Waymo has a strong R&D focus, actively publishing its latest research findings. Aurora company focuses on autonomous solutions for trucks serving logistics purposes. Its prediction module [^3] https://blog.aurora.tech/ performs joint object detection and trajectory forecasting, combining learning-based methods with rule-based constraints to align predictions with driving rules. To ensure safety across a wide range of scenarios, from highway driving delivering cargo between cities to urban logistics, Aurora’s system is tested extensively through a combination of virtual testing suites and real-world operations. Each company’s strategy reflects a balance between predictive accuracy and operational constraints, tailored to the specific purpose of their vehicles.
@@
For evaluation and testing, two robust open-source platforms, Autoware [^4] https://github.com/autowarefoundation/autoware and Apollo [^5] https://github.com/ApolloAuto/apollo, provide environments to test perception modules. Autoware’s perception stack includes functionalities such as object recognition (2D/3D detection using LiDAR, cameras, and radar), obstacle segmentation (distinguishing drivable areas via point cloud processing), occupancy grid mapping (representing occupied and free spaces), and traffic light recognition (detecting light states and directions). However, it lacks an integrated prediction module. Starting from release 5.0, Apollo incorporates an integrated prediction module [^6] https://github.com/ApolloAuto/apollo/blob/master/modules/prediction/README.md, supporting a range of predictors: Free Move Predictor for non-lane-constrained obstacles, Single Lane Predictor for vehicles on a single lane, and Lane Sequence Predictor for multi-lane and behavior-based trajectory predictions. Later versions, such as 6.0.0 and 7.0.0, introduced advanced prediction models, including a semantic map-based pedestrian prediction model and the Inter-TNT prediction model, which leverage interactive prediction and planning evaluation. While Apollo provides a comprehensive platform for testing prediction algorithms within the full autonomous driving stack, there is a clear need for an open, lightweight platform that enables faster experimentation with prediction modules. Such a platform should integrate a perception component to provide realistic inputs for testing, while remaining flexible enough to support subsequent integration into full-stack systems.
@@
For researchers entering the field of trajectory prediction, the volume of conducted works and proposed methods might be overwhelming, raising questions about how to categorize and classify existing approaches and which areas still offer room for improvement. For professionals experienced in the realm, keeping track of new emerging designs that lead to significant performance improvements, considering how rapidly the field is evolving, might be challenging. To assist both groups, this review covers a wide range of topics relevant to the prediction field, provides comprehensive overviews, and lists relevant references for readers to fetch additional information, which is either out of scope or briefly mentioned in the paper. In this review, we pose the following research questions:
@@
    RQ2: What are the emerging research trends in the trajectory forecasting realm? This question aims to discern the current focus of researchers and which pressing challenges they are addressing.
@@
    RQ3: Is trajectory prediction problem solved or there is a room for improvement? To answer this question, we discuss the effectiveness of current state-of-the-art (SOTA) prediction modules are, what are current limitations and which research gaps are still existing.
@@
    Provide an extensive overview of the prediction pipeline, which includes discussions of input modalities, modeling features and approaches, output modalities, evaluation metrics, and prediction paradigms.
@@
    Devise a taxonomy and categorization of modeling approaches, offering a structured classification of methods. Each category includes a comparative table summarizing key models and their characteristics for quick reference.
@@
    Provide an exhaustive literature review of active research areas in the trajectory prediction domain, covering topics such as uncertainty estimation, integration of driving knowledge, planning-conditioned prediction, vision-language model (VLM)-guided prediction models, and collaboration between vehicles.
@@
    Provide an open repository [^7] https://github.com/AV-Lab/Trajectory_Prediction_Models_Zoo featuring two prediction pipelines to facilitate experimentation, benchmarking, and the development of new methods. The first pipeline includes multiple learning-based models that can be trained on SOTA datasets or custom datasets, enabling flexibility for diverse research needs. The second pipeline implements an integrated detect-track-predict approach, designed to process inputs from specified sensor configurations (e.g., frontal camera, stereo camera, LiDAR) for seamless trajectory forecasting in real-world scenarios.
@@ 2 Overview of Prediction Pipeline
This section aims to provide a comprehensive overview of a prediction pipeline. Figure 3 illustrates the general prediction pipeline, highlighting its main components. The overview begins with the problem formulation of trajectory prediction task, elaborates on input modalities, provides a discussion of modelling features and approaches, along with the classification of output modalities. The section continues by elaborating on metrics for performance evaluation and listing benchmark datasets with leader-board solutions. It concludes with a discussion of trajectory prediction paradigms.
@@ 2.2 Input Modalities
In this subsection, we discuss the classification of input modalities and review studies in the literature for each category. We classify input modalities into three groups: past trajectories of agents, sensor data, and HD maps.
@@ 2.2.1 Past Trajectories
Past trajectory is a fundamental input in trajectory prediction models. Predicting future trajectories solely based on past trajectories [17, 18, 19, 20] forces the models to learn patterns by relying purely on spatio-temporal dependencies. This approach, which excludes additional contextual or environmental information, is efficient in terms of prediction speed and computational cost. However, it lacks the capacity to capture knowledge about the environment or interactions between agents, making it less robust in complex scenarios. One way to implicitly encode interactions is to include the past trajectories of surrounding agents in the prediction process. Studies [21, 22, 23, 24, 25, 26, 27] have demonstrated that including the trajectories of neighboring agents allows the model to infer interdependencies and interactions among them. This enables predictions that are more reflective of real-world behaviors, such as reactions to nearby vehicles or pedestrians. However, the main limitation of these models is reliance on error-free observations during training, using ground-truth past trajectories. In practical applications, the past trajectories available to prediction models are often noisy or incomplete (when we just start tracking the object, the missing observations will be zeros, until we accumulate the past trajectory of the required length), and robust handling of these inaccuracies is critical for practical deployment. When discussing past trajectories, they are typically represented as $(x,y)$ coordinates over multiple timestamps. However, additional motion states, such as heading and speed, can complement this information. These inputs can either be included as part of the agent’s state across multiple past timestamps or represent characteristics of the agent’s current state. Studies [28, 29, 30] have highlighted the importance of including velocity, acceleration, and orientation as input features to enhance prediction accuracy.
Models relying solely on past trajectories assume error-free perception, as they are trained on ground-truth data. In reality, perception systems often produce noisy or incomplete inputs. To address this, either the perception pipeline must be highly robust, or prediction models should incorporate additional input modalities, from which environmental context or agent interactions can be extracted, to improve accuracy and compensate for perception errors.
@@ 2.2.2 Sensors Data
Sensor data, provided by exteroceptive sensors such as LiDAR, radar, and cameras, enables prediction models to capture environmental context and integrate it into predictions. This section first examines each sensor individually and then explores approaches that leverage multi-modal inputs for enhanced prediction performance.
@@
Two significant limitations hinder the widespread adoption of LiDAR technology are its relatively high cost and its accuracy in detecting pedestrians. Although LiDAR technology is advancing and prices are gradually declining, it remains expensive and impractical for consumer-level mass-market adoption. The second limitation is its lower accuracy in pedestrian detection. For example, on the KITTI leaderboard, the top pedestrian detection score is only 56% [42], compared to a 93% vehicle detection score achieved by [43]. This disparity stems from LiDAR’s lack of rich semantic information, which is crucial for accurate pedestrian detection. As a result, LiDAR systems are ideally complemented by cameras, which provide the necessary semantic context to enhance pedestrian detection performance.
Camera. Cameras play a crucial role in perception by scanning the environment and detecting traffic participants, road signs, and traffic lights. These sensors are characterized by their resolution, field of view, and configuration. RGB cameras are the standard visual sensors for AVs and are widely used to collect data for current SOTA datasets. The vision community for RGB videos is highly advanced, providing efficient algorithms for object detection [44], semantic segmentation [45], and lane recognition [46]. Camera setups in AV systems are generally categorized as single frontal cameras [47] or multi-view configurations [48]. Single frontal camera setups primarily rely on 2D perception and are mostly used in driving assistance systems. In contrast, multi-view setups arrange cameras to ensure complementary, non-overlapping fields of view, providing 360-degree observability. This configuration enables 3D perception [49] and close-range monitoring, especially in urban and dynamic driving conditions. Additionally, camera systems can employ either monocular or stereo setups. Monocular cameras [44, 45, 46, 49] are lightweight and cost-effective but restricted to 2D image data, relying on advanced algorithms for depth estimation. In contrast, stereo cameras utilize dual lenses to calculate depth through stereo triangulation, enabling accurate 3D perception [50], which is essential for tasks like depth estimation and obstacle avoidance.
@@
Kim et al. [56] presented a model to classify agent types (pedestrian, cyclist, or vehicle) using radar data. Radar signals are transformed into range-velocity (RV) images that capture an object’s distance and motion. These RV images are processed by a convolutional LSTM to learn spatial and temporal features, followed by a fully connected layer to map these features to target classes. Steinhauser et al. [57] proposed a system focused on individual limb movements for pedestrian behavior prediction. Their method employs a deconvolution algorithm combined with a clustering technique to extract and analyze reflections from pedestrian body parts, decomposing the range-Doppler spectrum to identify the torso, arms, and legs. Despite all advancements, radar remains primarily a supplementary sensor in SOTA models, complementing cameras and LiDAR by providing reliable velocity and heading information.
Rasterization. Rasterization converts raw sensor data into structured, grid-like representations, enabling prediction models to effectively process spatial and temporal information. For example, LiDAR point clouds can be rasterized into occupancy grids, where each cell indicates whether it is occupied or free. Radar data adds dynamic attributes such as velocity and direction [58], which can be encoded either through color coding (e.g., higher speeds represented by brighter colors) [59] or as separate velocity channels, storing velocity values in an additional data layer for each grid cell. Similarly, camera images can be rasterized to map rich semantic information, such as object classes and bounding boxes. This unified rasterized format facilitates the fusion of multimodal sensor data and enables machine learning models to process environmental context with high precision.
The most common type of rasterization in autonomous vehicles is Bird’s Eye View (BEV) rasterization [60, 61, 62, 63, 64, 65], where sensor data is transformed into a top-down perspective of the environment. In BEV rasterization, dynamic parameters such as positions and velocities are mapped into a unified plane aligned with the ego vehicle. This format often employs color-coded semantics to differentiate objects and incorporates temporal information through stacked frames or brightness variations to represent motion over time. Simplified BEV representations extend this concept by depicting static and dynamic elements, e.g., road lanes, vehicles, and pedestrians, using polygons and lines.

### v2

@@ Abstract
As the potential for autonomous vehicles to be integrated on a large scale into modern traffic systems continues to grow, ensuring safe navigation in dynamic environments is crucial for smooth integration. To guarantee safety and prevent collisions, autonomous vehicles must be capable of accurately predicting the trajectories of surrounding traffic agents. Over the past decade, significant efforts from both academia and industry have been dedicated to designing solutions for precise trajectory forecasting. These efforts have produced a diverse range of approaches, raising questions about the differences between these methods and whether trajectory prediction challenges have been fully addressed. This paper reviews a substantial portion of recent trajectory prediction methods proposing a taxonomy to classify existing solutions. A general overview of the prediction pipeline is also provided, covering input and output modalities, modeling features, and prediction paradigms existing in the literature. In addition, the paper discusses active research areas within trajectory prediction, addresses the posed research questions, and highlights the remaining research gaps and challenges.
@@ 1 Introduction
To accurately predict trajectories, the prediction model must account for the motion of vehicles, cyclists, pedestrians — all traffic participants. For vehicle motion, autonomous vehicles (AV) should be able to track surrounding vehicles and their potential maneuvers. As human drivers have an intuition of inter-dependency of surrounding vehicles, e.g., it can be anticipated if a neighboring car is cut off, there is a high chance that this vehicle will be compelled to perform a risky maneuver and potentially interfere with the ego-vehicle’s lane to avoid collision or suddenly hit the breaks. To equip autonomous vehicles with interaction-aware intuition, the prediction module needs to build a representation of each target agent’s surroundings and accounting for inter-dependency of traffic participants’ decisions. Regarding the human motion, beyond the complexity of interaction and inter-dependency, predicting human trajectories involves accounting for complex human behavior. Pedestrians’ motion is influenced not only by road topology and actions of surrounding agents but also by social rules, norms, and relations between agents. To accurately predict pedestrian trajectories, many factors need to be inferred from contextual and general knowledge information, e.g., a child and an adult exhibit different motion patterns. Additionally, prediction models that rely on traffic rules benefit from a reduced set of possible future trajectories. However, human drivers often violate traffic rules by over-speeding, changing lanes through a solid line, turning from a non-designated lane, ignoring stop signs, or even passing through a red light. Prediction models need to account for the risk of such violations to ensure robust and reliable predictions in real-world scenarios. Lastly, the trajectory prediction task has a multimodal nature, i.e., the same past trajectory can unfold into several possible future trajectories, e.g., a vehicle approaching an intersection in a lane marked for both straight and left-turn movements may continue forward or make a left turn. To account for multimodality, the output of the prediction model must contain the distribution of possible trajectories with probabilities.
@@
To address these challenges, the research community has focused on designing models capable of capturing complex patterns essential for accurate prediction. Among the top-performing solutions is Trajectron++ [1], which integrates environmental context through a graph-structured recurrent architecture and Spatio-Temporal Transformer Network (S2TNet) [2] leveraging transformer layers to capture both spatial and temporal dependencies. Further, from the intersection between academia and industry, the MultiPath++ model [3], released by the Waymo research group, employs multi-context gating fusion and dynamically learns reference points (anchors) to support goal-oriented predictions. Collectively, these models combine diverse techniques, such as graph-based interaction modeling, transformer architectures, and dynamic goal sampling, leading to more accurate and robust trajectory predictions. On the other side, the commercial research prioritizes real-time execution, cost-effectiveness, and handling uncertain situations. This leads each company to use its own combination of sensors, compatible with the set of algorithms it employs. Tesla’s prediction algorithm is an end-to-end vision-based deep learning model [^1] https://www.tesla.com/support/autopilot, forecasting trajectories based on the visual sequences and aggregated contextual information over time. The model continuously learns from the vast amount of data collected by Tesla’s global fleet of vehicles. In contrast, Waymo utilizes a combination of high-fidelity sensors (LiDAR, cameras, and radar). The sensor data is processed by the perception module and then passed to motion prediction and planning components. Waymo adopts a modular structure, ensuring that each component can be thoroughly tested independently and then integrated into a single pipeline [^2] https://waymo.com/research/. It is challenging to determine the specific algorithms employed in real operations, but Waymo has a strong R&D focus, actively publishing its latest research findings. Aurora company focuses on autonomous solutions for trucks serving logistics purposes. Its prediction module [^3] https://blog.aurora.tech/ performs joint object detection and trajectory forecasting, combining learning-based methods with rule-based constraints to align predictions with driving rules. To ensure safety across a wide range of scenarios, from highway driving delivering cargo between cities to urban logistics, Aurora’s system is tested extensively through a combination of virtual testing suites and real-world operations. Each company’s strategy reflects a balance between predictive accuracy and operational constraints, tailored to the specific purpose of their vehicles.
@@
On a testing side of autonomous driving algorithms, two robust open-source platforms provide environment for evaluation: Autoware [^4] https://github.com/autowarefoundation/autoware and Apollo [^5] https://github.com/ApolloAuto/apollo. Autoware’s perception stack includes functionalities such as object recognition (2D/3D detection using LiDAR, cameras, and radar), obstacle segmentation (distinguishing drivable areas via point cloud processing), occupancy grid mapping (representing occupied and free spaces), and traffic light recognition (detecting light states and directions). However, it lacks an integrated prediction module. Starting from release 5.0, Apollo incorporates an integrated prediction module [^6] https://github.com/ApolloAuto/apollo/blob/master/modules/prediction/README.md, supporting a range of predictors: Free Move Predictor for non-lane-constrained obstacles, Single Lane Predictor for vehicles on a single lane, and Lane Sequence Predictor for multi-lane and behavior-based trajectory predictions. Later versions, such as 6.0.0 and 7.0.0, introduced advanced prediction models, including a semantic map-based pedestrian prediction model and the Inter-TNT prediction model with interactive prediction and planning evaluation. While Apollo provides a comprehensive platform for testing prediction algorithms within the full autonomous driving stack, there is a clear need for an open, lightweight platform that enables faster experimentation with prediction modules. Such a platform should integrate a perception component to provide realistic inputs for testing, while remaining flexible enough to support subsequent integration into full-stack systems.
@@
For researchers entering the field of trajectory prediction, the volume of conducted works and proposed methods might be overwhelming, raising questions about how to categorize and classify existing approaches and which areas still offer room for improvement. For professionals experienced in the realm, keeping track of new emerging designs that lead to significant performance improvements, considering how rapidly the field is evolving, might be challenging. To assist both groups, this review covers a wide range of topics relevant to the trajectory prediction field, provides comprehensive overviews, and lists relevant references for readers to fetch additional information, which is either out of scope or briefly mentioned in the paper. In this review, we pose the following research questions:
@@
    RQ2: Is trajectory prediction problem solved or there is a room for improvement? To answer this question, we reflect on the current state of trajectory prediction research, outlining the challenges and research gaps guiding ongoing development, as well as the limitations that hinder real-world deployment and integration.
@@
    RQ3: What are the emerging research trends in the trajectory forecasting realm? This question aims to discern the current focus of researchers and which pressing challenges they are addressing.
@@
    Provide an extensive overview of the prediction pipeline covering input modalities, modeling features and approaches, output modalities, evaluation metrics, and prediction paradigms.
@@
    Extend the taxonomies presented in [4, 5], offering a structured classification of modeling approaches and positioning the broad set of methods found in the literature. Each category is supported by a comparative table summarizing key models and their defining characteristics for quick reference.
@@
    Provide a review of active research areas in the trajectory prediction domain, covering topics such as uncertainty estimation, integration of driving knowledge, planning-conditioned prediction, vision-language model (VLM)-guided prediction models, and collaboration between vehicles.
@@
    Provide an open repository [^7] https://github.com/AV-Lab/Trajectory_Prediction_Models_Zoo featuring two prediction pipelines to facilitate experimentation, benchmarking, and the development of new methods. The first pipeline includes multiple learning-based models that can be trained on SOTA or custom datasets, enabling flexibility for diverse research needs. The second pipeline implements a detect–track–predict approach that processes inputs from various sensor configurations (e.g., frontal camera, stereo camera, LiDAR), enabling emulation of real-world deployment scenarios.
@@ 2 Overview of Prediction Pipeline
This section aims to provide a comprehensive overview of a prediction pipeline. Figure 3 illustrates the general prediction pipeline flow, highlighting its main components. The overview begins with the problem formulation of trajectory prediction task, elaborates on input modalities, provides a discussion of modelling features and approaches, and iterates through possible output modalities. The section continues by elaborating on metrics used to evaluate trajectory forecasting models and listing benchmark datasets with leader-board solutions. It concludes with a discussion of existing in the literature trajectory prediction paradigms.
@@ 2.2 Input Modalities
In this subsection, we discuss the classification of input modalities and review relevant studies in the literature for each category. Input modalities are classified into three groups: past trajectories of agents, sensor data from LiDAR, radar, or cameras, and high-definition (HD) maps describing static environmental context. Explicitly encoding tracked past trajectories is one way to capture temporal dynamics; alternatively, temporal information can be inferred from sequences of raw sensor inputs. Regardless of the format, providing input from which dynamics can be extracted is a fundamental cornerstone of trajectory prediction models.
The combination of input modalities, e.g., coupling past trajectories with HD maps for road layout information and camera feeds for traffic light detection, is shaped by the system’s sensing capabilities, operational constraints, and design choices. Models trained solely on past trajectories may require complementary context under noisy conditions, which can be addressed by processing additional sensor data to extract semantic and environmental cues. While HD maps enrich predictions with high-level semantics, they also introduce dependencies on precise localization and map availability. Each modality has inherent limitations, and combining them increases processing complexity. In the following subsections, we discuss each input modality in detail, highlighting its strengths and limitations.
@@ 2.2.1 Past Trajectories
Past trajectory is a fundamental input in trajectory prediction models. Predicting future trajectories solely based on past trajectories [17, 18, 19, 20] forces the models to learn patterns by relying purely on spatio-temporal dependencies. This approach, which excludes additional contextual or environmental information, is efficient in terms of prediction speed and computational cost. However, it lacks the capacity to capture knowledge about the environment or interactions between agents, making it less robust in complex scenarios. One way to implicitly encode interaction of the target agent is to include the past trajectories of its surrounding agents in the prediction process, as shown in the studies [21, 22, 23, 24, 25, 26, 27]. This enables predictions that are more reflective of real-world behaviors, such as reactions to nearby vehicles or pedestrians. However, the main limitation of these models is reliance on one input and assumption of error-free observations. The models are trained using ground-truth past trajectories, whereas in practical applications, the past trajectories available to prediction models are often noisy or incomplete, and robust handling of these inaccuracies is critical for practical deployment. Typically the agent state in past trajectories is represented as $(x,y)$ coordinates. Additionally, heading and speed can be added complementary. Studies [28, 29, 30] have highlighted the importance of including velocity, acceleration, and orientation as input features to enhance prediction accuracy.
@@ 2.2.2 Sensors Data
Sensor data, provided by exteroceptive sensors such as LiDAR, radar, and cameras, enables prediction models to capture environmental context for trajectory prediction models. This section first examines each sensor individually and then discusses approaches of integrating multi-modal inputs.
@@
Two significant limitations hinder the widespread adoption of LiDAR technology: its relatively high cost and its accuracy in detecting pedestrians. Although LiDAR technology is advancing and prices are gradually declining, it remains expensive and impractical for consumer-level mass-market adoption. The second limitation is its lower accuracy in recognizing fine-grained classes such as pedestrians and cyclists. This disparity stems from its sparse, geometry-centric output and lack of rich semantic information, which makes interpretation of complex objects challenging. For example, on the KITTI leaderboard, the top pedestrian detection score is only 56% [42], compared to a 93% vehicle detection score achieved by [43]. As a result, LiDAR systems are ideally complemented by cameras, which provide the necessary semantic context to enhance pedestrian detection performance.
Camera. Cameras play a crucial role in perception of autonomous vehicles in detecting traffic participants, road signs, and traffic lights. These sensors are characterized by their resolution, field of view, and configuration. The vision community is highly advanced, providing efficient algorithms for object detection [44], semantic segmentation [45], and lane recognition [46]. Camera setups in AV systems are generally categorized as single frontal cameras [47] or multi-view configurations [48]. Single frontal camera setups primarily rely on 2D perception and are mostly used in driving assistance systems. In contrast, multi-view setups [49] arrange cameras to ensure complementary, non-overlapping fields of view, providing 360-degree observability. Additionally, camera systems employ either monocular or stereo setups. Monocular cameras [44, 45, 46, 49] are lightweight and cost-effective but restricted to 2D image data, relying on advanced algorithms for depth estimation. Stereo cameras utilize dual lenses to calculate depth through stereo triangulation, enabling 3D perception [50].
@@
Kim et al. [56] presented a model to classify agent types (pedestrian, cyclist, or vehicle) using radar data. Radar signals are transformed into range-velocity (RV) images that capture an object’s distance and motion. These RV images are processed by a convolutional LSTM to learn spatial and temporal features, followed by a fully connected layer to map these features to target classes. Steinhauser et al. [57] proposed a system focused on individual limb movements for pedestrian behavior prediction. Their method employs a deconvolution algorithm combined with a clustering technique to extract and analyze reflections from pedestrian body parts, decomposing the range-Doppler spectrum to identify the torso, arms, and legs. Radar typically complements LiDAR and camera inputs in SOTA datasets, providing reliable velocity and heading cues.
Rasterization. Rasterization converts raw sensor data into structured, grid-like representations, enabling prediction models to effectively process spatial and temporal information. For example, LiDAR point clouds are can be transformed into grid-based formats, such as intensity or height maps, that capture spatial geometry. Radar data adds dynamic attributes such as velocity and direction [58], which can be encoded either through color coding (e.g., higher speeds represented by brighter colors) [59] or as separate velocity channels, storing velocity values in an additional data layer for each grid cell. Similarly, camera images can be rasterized to map rich semantic information, such as object classes and bounding boxes. This unified rasterized format facilitates the fusion of multimodal sensor data.
The most common type of rasterization for autonomous driving is Bird’s Eye View (BEV) rasterization [60, 61, 62, 63, 64, 65], where sensor data is transformed into a top-down perspective of the environment. In BEV rasterization, dynamic parameters such as positions and velocities are mapped into a unified plane aligned with the ego vehicle. This format often employs color-coded semantics to differentiate objects and incorporates temporal information through stacked frames or brightness variations to represent motion over time. Simplified BEV representations extend this concept by depicting static and dynamic elements, e.g., road lanes, vehicles, and pedestrians, using polygons and lines.